# DR-MPC: Deep Residual Model Predictive Control for Real-world Social Navigation

### v1

@@ V The Pipeline
The exploration-exploitation problem in DRL does not have to be addressed solely by the actor network. To perform better than random exploration, we guide the selection of $\mathbf{a}^{i}_{HA}$ using heuristics to find promising directions for human avoidance. The sooner these valuable experiences are stored in the replay buffer, the sooner the critic can correct these Q-values. This concept of early intervention in uncertain or high-risk states has proven effective in interactive imitation learning and sample-efficient DRL [29, 30].
To implement this, we perform OOD state detection to assess the DRL model’s familiarity with a state. If the state is in-distribution, we execute DR-MPC’s action. Otherwise, we either validate the model’s action as safe using a heuristic safety check or override it using a heuristic policy. This process steers the DRL agent toward collision-free areas.
Once a state is in-distribution, the DRL agent will continue to explore and may still encounter collisions. The agent will quickly learn to focus on known good actions and avoid poor reward regions, thereby accelerating learning speed. Note that the OOD state detection, heuristic safety check, and heuristic policy are modular components and can be swapped out or fine-tuned independently of everything else in the pipeline.
@@ V-1 OOD State Detection
There are numerous methods for OOD state detection: we select a $K$-nearest neighbours (KNN) algorithm in the latent space of the model [31]. This method is relatively interpretable and can be implemented in real-time using Facebook’s Faiss package [32]. We extract the latent embedding from our model at the layer immediately following the fusion of path tracking and human avoidance information.
Figure 4: Simulation results averaged over 12 trials. Both DR-MPC models outperform naïve DRL and Residual DRL. While DRL and Residual DRL learn to achieve a high path advancement reward, they struggle with task switching to perform proficient human avoidance.
@@ V-2 Heuristic Safety Check and Policy
Figure 5: Simulation training scenarios. The path is black, the virutal corridor is red, the robot is yellow, and the humans are light grey.
Upon detecting an OOD state, we assess whether the DRL model’s proposed action is safe. We assume a constant velocity motion model (CVMM) for both the human and the robot. We roll out the proposed action for the robot and the estimated human velocity based on the last two timesteps for a predefined number of timesteps. We then evaluate if the safety-human raise is triggered at any point during this rollout; if it is, the action is deemed unsafe. For the safety-corridor condition, we only check the one-timestep rollout result.
@@ VI-A Simulator and Setup
We augment the CrowdNav simulator from [10], originally developed by [6]. In a modularized manner, we replace waypoint navigation with path tracking. Our training schema cycles through four scenarios, as illustrated in Figure 5. Once a path is completed, the next one begins. We run our simulator with $n_{t}=6$; these 6 humans are modeled using ORCA and continuously move to random goals in the arena. Our robot has an action space of $v\in[0,1]$ and $\omega\in[-1,1]$. The MDP has a timestep of 0.25s, and we train our models with a limited amount of data: 37,500 steps—around 2.5 hours of data.
@@ VI-B Model Baselines
We evaluate four models. The first applies the model of [10], which shares a similar backbone to DR-MPC. The DRL model combines the ‘PT Embedding’ and ‘HA Embedding’ with an MLP to generate the agent’s action; we refer to this model as the naïve DRL model. The second model is a residual DRL model guided by [20], which outputs a corrective action: the action executed in the environment is the sum of the MPC and DRL corrective action, which is then truncated into the feasible action space. Since path tracking and human avoidance may require distinct actions, we set the DRL corrective action to have $v\in[-1,1]$ and $\omega\in[-2,2]$, ensuring that the final action can cover the entire action space. The residual DRL model architecture is the same as the naïve DRL model but also has $\mathbf{a}_{\text{MPC}}$ inputted into the final MLP. Finally, we evaluate two versions of DR-MPC: one without the OOD state detection and CVMM modules, and one with them.
@@ VI-C Results and Discussion
Figure 6: Real-world demonstrations. Yellow is the reference path, orange is the robot’s trajectory, and purple is the human’s trajectory. Left and middle: the robot learns to navigate around humans standing in the robot’s path. Right: The robot learns to deviate from its path to avoid disturbing the human’s trajectory.
The main results are depicted in Figure 4. The plotted lines are the mean values across 12 trials, while the shaded areas indicate the variance observed during training. For all metrics, the values in these plots are averaged over 500 environment steps. We do not use the reward-per-episode metric because different path lengths inherently yield different maximum cumulative rewards, which would skew comparisons.
From the cumulative reward plot, it is clear that both of our DR-MPC models significantly outperform the naïve DRL and residual DRL models. While all models can learn a high path-advancement reward and navigate paths effectively, interacting with humans remains a greater challenge. DR-MPC, however, excels in task switching and optimizing both path tracking and human avoidance using the $\boldsymbol{\alpha}$ parameter. It is important to note that the DRL and residual DRL models were verified for correctness and, given sufficient training, could eventually reach similar reward levels as DR-MPC.
@@
Comparing our DR-MPC models, the one with OOD state detection performs better in the early stages of training. This advantage comes from the heuristic policy exploring actions deemed safe by the CVMM safety check and that are close to the MPC action, which results in higher path-advancement rewards. The similarity in safety interventions between the two models suggests that ORCA, the human motion model, is not well captured by a CVMM in close contact scenarios. However, the CVMM safety check is valuable in real-world situations, particularly with inattentive humans blindly walking on a straight path or humans who are standing still.
Overall, the results demonstrate that the DR-MPC approach seamlessly integrates MPC behaviour and adjusts to dynamic environments making DR-MPC a robust and efficient solution.
@@ VII-A Implementation
VII-A1 Robot
We use a Clearpath Robotics Jackal equipped with an Ouster OS0-128 LiDAR. The OS0 sensor has a $90^{\circ}$ vertical field-of-view (FOV) and a $360^{\circ}$ horizontal FOV, capturing range and reflectivity data. This data can be represented as equirectangular images—an example of the reflectivity image is shown in Figure 2. These images are comparable to low-resolution cameras, enabling the application of computer-vision models. By fusing the 2D computer-vision results with the LiDAR’s depth information, we can produce 3D outputs.
VII-A2 Mapping, Localization, Path Tracking
We use the open-source Teach and Repeat (T&R) codebase [34]. Specifically, in LiDAR T&R, after manually driving the robot through a new environment once, the robot can subsequently localize itself to this previously driven path and track it using MPC. T&R enables rapid deployment of our pipeline in new environments.
@@ VII-A2 Mapping, Localization, Path Tracking
VII-A3 Human Tracking
We detect humans by running a pre-trained YOLOX model from mmDetection on the reflectivity image and its $180^{\circ}$ shifted version to account for humans at the wrap-around point [35]. By combining depth information from the range image with the localization result from T&R, we recover the 3D world positions of humans in the scene and track them using the Norfair package [36].
@@ VII-B Experimental Setup
Although we trained on 3.75 hours of experience transitions, the entire process took about 15 hours. After every 500 experience tuples, we paused data collection to allow the model to update, performing twice as many training updates as environment samples. While off-policy learning can leverage older data, keeping the inference model synchronized with the training model allows Q-values to be corrected as state-action distributions shift. Additionally, outliers caused by processing delays from resource allocation variance were filtered to reduce data noise. Also, soft resets, along with charging the laptop and robot batteries, contributed to the additional time.
@@
The real world introduces challenges absent in simulation, such as perception errors (missed detections, false positives, misidentifications) and variable timing delays from tasks (localization, human detection, action generation). Also, human movements are probabilistic, and the robot faces acceleration constraints based on battery charge. These challenges highlight a key advantage of DRL: it maximizes expected cumulative reward, allowing effective learning in noisy environments.
@@ VII-C Results and Discussion
TABLE I: Real-world Comparison of Final Policies ($\mu\pm\sigma$ Per Loop)
Algorithm Steps to Goal ($\downarrow$) Safety-human Safety-corridor
Raises ($\downarrow$) Raises ($\downarrow$)
@@
Table I shows that the final DR-MPC model outperforms benchmark models in all key metrics—fewer safety raises and faster navigation. The training process (Figure 7) begins with performance nearly identical to the heuristic policy (as expected). During exploration, DR-MPC performs worse but eventually coalesces its experience to achieve superior results. After 4 hours of data, DR-MPC’s performance begins to plateau. While more data would add marginal improvements, it is unlikely to completely eliminate safety-human raises. Social navigation is a long-tail problem with rare events, and even in simulation with a deterministic agent, safety-human raises never reach zero.
During training, $\beta$ (orange) trends toward 0, while the average $\alpha$ (blue) remains low, indicating the model learns when to best select the MPC action. There is also an upward trend for in-distribution (green) states. The spike at 5000 steps is not an artifact—early training has major changes in latent representations, making rare state identification difficult, but it quickly resolves with a few thousand environment steps.
Qualitatively, DR-MPC exhibits key crowd-navigation behaviors: smoothly weaving around still humans (left and middle of Figure 6) and deviating from its path and slowing down for moving humans (Figure 1 and right of Figure 6). In human-free areas, DR-MPC switches to MPC path tracking.
@@ Appendix A Human Avoidance Network
For each human trajectory $\mathbf{q}^{t-H_{i}:t}_{i}=\begin{bmatrix}\mathbf{p}^{t-H_{i}}_{i}\dots\mathbf{p}^{t}_{i}\end{bmatrix}$, we sequentially process it from time $t-H_{i}$ to $t$ using a GRU to generate an embedded trajectory $\mathbf{e}^{i}_{\text{traj}}$. This step handles human trajectories of varying lengths, embedding them into a uniform latent space. Next, with $\mathbf{E}_{\text{traj}}=\begin{bmatrix}\mathbf{e}^{1}_{\text{traj}}\dots\mathbf{e}^{n_{t}}_{\text{traj}}\end{bmatrix}$, we use three MLPs to generate the queries $\mathbf{Q}_{\text{traj}}$, keys $\mathbf{K}_{\text{traj}}$, and values $\mathbf{V}_{\text{traj}}$. We then apply multi-head attention using the scaled dot-product attention: $\text{MultiHead}(\mathbf{Q}_{\text{traj}},\mathbf{K}_{\text{traj}},\mathbf{V}_{\text{traj}})$ (see [37]). The output of this module is $\mathbf{E}_{\text{HH}}$, a $n_{t}\times d_{HH}$ tensor, where $d_{HH}$ is the dimension of the human-human embeddings.

### v2

@@ V The Pipeline
To perform better than random exploration, we guide the selection of $\mathbf{a}^{i}_{HA}$. Early intervention in uncertain or high-risk states has proven effective in interactive imitation learning [29]. We perform OOD state detection: if the state is in-distribution (ID), we execute DR-MPC’s action. Otherwise, we either validate the model’s action as safe using a heuristic safety check or override it using a heuristic policy to steer the robot toward collision-free areas.
Once a state is ID, the DRL agent will continue to explore and may still encounter collisions. The agent will quickly learn to focus on known good actions and avoid poor reward regions, thereby accelerating learning speed. Note that the OOD state detection, heuristic safety check, and heuristic policy are modular components and can be swapped out or fine-tuned independently of everything else in the pipeline.
@@ V-1 OOD State Detection
We use the SOTA OOD algorithm: $K$-nearest neighbors (KNN) in the latent space of our model [30]. We extract the model’s latent embedding for each state in the replay buffer. With the Faiss package [31], we can efficiently query the $K$’th closest vector. A state is considered OOD if its distance to the $K$’th nearest vector exceeds the threshold: the average distance between ($\mathcal{S}$, $\mathcal{S}^{\prime}$) pairs in the replay buffer. Intuitively, if a query state is ID, it will roughly have $K$ other states nearby. This method offers two key advantages over approaches like Random Network Distillation (RND) [32]. First, it requires minimal additional compute as it leverages the existing model’s latent space without needing to train a new model. Second, by explicitly utilizing all the data, it avoids capacity issues where methods such as RND can ‘forget’ due to the model’s limited capacity.
@@ V-2 Heuristic Safety Check and Policy
Figure 4: Simulation scenarios. The path is black, the corridors are red, the robot is yellow, and the humans are grey.
Upon detecting an OOD state, we assess the safety of DR-MPC’s action using a constant velocity motion model (CVMM). We roll out the robot’s proposed action and the human’s estimated motion (position difference in the last two timesteps) for two seconds. If a safety-human raise is triggered in the rollout, the action is deemed unsafe. For the safety-corridor condition, we evaluate the one-timestep rollout result.
@@ VI-A Simulator and Setup
We augment the CrowdNav simulator from [10] by replacing waypoint navigation with path tracking. Our training schema cycles through four scenarios (Figure 4). We run our simulator with $n_{t}=6$; these 6 humans are modeled using ORCA and continuously move to random goals in the arena. As in [19], because we use a disturbance penalty, the robot is visible to the humans. Our robot has an action space of $v\in[0,1]$ and $\omega\in[-1,1]$. The MDP has a timestep of 0.25s, and we train our models with a limited amount of data: 37,500 steps—around 2.5 hours of data. We set $K=100$ so that by the end of training, over $95\%$ of the states are ID.
@@ VI-B Model Baselines
We evaluate five models. The first applies [10], which shares a similar backbone to DR-MPC. This model combines the ‘PT Embedding’ and ‘HA Embedding’ with an MLP to generate the agent’s action; we refer to this model as the naïve DRL model. The second model is residual DRL [20], which outputs a corrective action: the action executed in the environment is the sum of the MPC and DRL action truncated into the feasible action space. The corrective action has range $v\in[-1,1]$ and $\omega\in[-2,2]$ to ensure the summed action can cover the entire action space. The residual DRL model architecture is the same as the naïve DRL model but also has $\mathbf{a}_{\text{MPC}}$ inputted into the final MLP. We evaluate two versions of DR-MPC: one without the OOD state detection and CVMM modules, and one with them. Finally, we compare against ORCA, which, although it does not optimize for the same objectives as our DRL-based models, serves as an intuitive performance baseline.
@@ VI-C Results and Discussion
Figure 5 depicts training progress and Table I is the final model comparison. From the cumulative reward plot, both DR-MPC models significantly outperform the naïve DRL and Residual DRL models, which have difficulty advancing on the path while avoiding safety-human raises. DR-MPC, however, excels in task switching and optimizes both path tracking and human avoidance using the $\boldsymbol{\alpha}$ parameter. ORCA guarantees zero collisions with humans modeled as ORCA and achieves the highest success rate but with the highest NNT. Furthermore, ORCA does not attempt to minimize human comfort (disturbance) or deviation from the path, which is practically useful in the real world where staying closer to the desired path typically aligns better with drivable areas.
@@
Figure 5: Simulation results averaged over 10 trials. Both DR-MPC models outperform naïve DRL and Residual DRL by efficient task switching for human avoidance.
TABLE I: Performance Comparison. ‘SR’ is success rate, ‘EDR’ is end deviation rate, ‘SHRR’ is safety-human raise rate, ‘SCRR’ is safety-corridor raise rate, ‘ATR’ is actuation termination rate, and ‘NNT’ is normalized navigation time which is navigation time divided by path length.
[…]
ORCA 0.60 0.07 0.00* 0.33 0.00 1.34
DRL [10] 0.20 0.02 0.77 0.01 0.00 1.16
Residual 0.21 0.00 0.79 0.00 0.00 1.06
DRL [20]
DR-MPC 0.57 0.02 0.30 0.00 0.11 1.32
DR-MPC 0.58 0.06 0.31 0.01 0.04 1.17
w/ OOD
Comparing our DR-MPC models, the one with OOD state detection demonstrates better performance in the early stages of training, as the heuristic policy helps guide it away from collisions. As training progresses and more states become ID, we observe a temporary increase in safety-human raises and a dip in cumulative reward. Shortly after, its performance converges with the model without OOD detection. Thus, OOD detection enables greater initial performance while ultimately achieving similar long-term results.
We qualitatively analyze OOD detection by deploying our trained models in an environment with two additional static humans placed directly on the path. When the robot approaches a static human, OOD detection is triggered, and the heuristic policy guides the robot around the human. In the real-world results (Figure 6), the upward trend of the ID (green) line has a spike at 5000 steps, which is not an artifact but reflects early training challenges in latent representation changes making rare state identification difficult. To overcome this, we start with a stricter KNN distance threshold (we scale the threshold calculation by $\frac{1}{3}$) and gradually relax this scaling factor to identity by the end of training. This tuning leads to better results than a constant distance threshold.
Performance could be further improved with a better heuristic safety check and heuristic policy. Currently, the CVMM safety check is overly conservative: early in training, 21% of CVMM triggers result in DR-MPC collisions on the next timestep, dropping to 5% by the end of training. Similarly, early in training when the heuristic policy (which relies on CVMM) causes a collision, DR-MPC also collides 90% of the time, but this decreases to 3% by the end of training. These findings suggest that CVMM struggles to model close-contact ORCA behavior. However, the CVMM safety check remains valuable in real-world scenarios, particularly for inattentive humans walking straight or standing still.
VI-D Reward Ablation
Given our new decision process formulation, we analyze our reward function. Table II compares the base ‘DR-MPC w/ OOD’ model with all rewards to versions of the model where specific rewards are removed. Interestingly, without $r_{\text{pa}}$, DR-MPC still achieves a high SR, as the policy begins with MPC path-tracking behavior, allowing it to reach the end of the path; however, this results in a significantly higher NNT. Removing $r_{\text{dev}}$ increases EDR and SCRR as expected, while removing the goal penalty marginally increases EDR. Without $r_{\text{cor-col}}^{*}$, SHRR increases as expected. Removing $r_{\text{act}}^{*}$ results in similar overall performance but with higher ATR. Removing $r_{\text{hum-col}}^{*}$ causes SHRR to increase drastically. Lastly, removing $r_{\text{dist}}$ reduces SHRR and greatly increases episodes reaching the path’s end (SR + EDR). This result makes intuitive sense because $r_{\text{dist}}$ influences comfort rather than directly impacting the quantitative navigation metrics. Consequently, while its numerical benefits are evident, the qualitative impact it provides must also be carefully considered. Therefore, we conclude that all rewards are essential to our decision process formulation for real-world social navigation.
TABLE II: Reward Ablation Study on DR-MPC w/ OOD
[…]
Base 0.58 0.06 0.31 0.01 0.04 1.17
$r_{\text{pa}}$ 0.46 0.02 0.42 0.00 0.09 1.60
$r_{\text{dev}}$ 0.46 0.08 0.19 0.06 0.21 1.27
$r_{\text{goal}}^{*}$ 0.56 0.07 0.37 0.01 0.00 1.17
$r_{\text{cor-col}}^{*}$ 0.56 0.06 0.27 0.05 0.05 1.24
$r_{\text{act}}^{*}$ 0.56 0.09 0.24 0.04 0.08 1.30
$r_{\text{hum-col}}^{*}$ 0.24 0.00 0.76 0.00 0.00 1.07
$r_{\text{dist}}$ 0.59 0.13 0.21 0.00 0.07 1.54
@@ VII-A Implementation
We use a Clearpath Robotics Jackal equipped with an Ouster OS0-128 LiDAR that captures $360^{\circ}$ range and reflectivity data. This data can be represented as equirectangular images—an example of the reflectivity image is shown in Figure 2. These images are comparable to low-resolution cameras, enabling the application of computer-vision models. By fusing the 2D computer-vision results with the LiDAR’s depth information, we can produce 3D outputs.
For path tracking, we use Teach and Repeat (T&R) for rapid deployment in new environments [34]. Specifically, in LiDAR T&R, after manually driving the robot through a new environment once, the robot can subsequently localize itself to this previously driven path and track it using MPC.
@@
Figure 8: Real-world examples. Reference path is yellow, robot’s trajectory is orange, and human’s trajectory is purple. Left: robot navigates around two static humans. Right: robot deviates from its path to avoid disturbing the human.
Lastly, we detect humans by running a pre-trained YOLOX model on the reflectivity image and its $180^{\circ}$ shifted version to account for the wrap-around point [35]. By combining depth information with the localization result from T&R, we recover the 3D world positions of humans and track them using the Norfair package [36].
@@ VII-B Experimental Setup
Although we trained on 3.75 hours of experience transitions, the entire process took about 15 hours. After every 500 experience tuples, we paused data collection to allow the model to update, performing twice as many training updates as environment samples. Additionally, outliers caused by processing delays from resource allocation variance were filtered to reduce data noise. Also, soft resets, along with charging the laptop and robot batteries, contributed to the additional time.
@@
The real world introduces challenges absent in simulation, such as perception errors (missed and false detections), variable processing delays (localization, human detection, action generation), probabilistic human motion, and robot acceleration dependent on battery charge. These challenges highlight a key advantage of DRL: it maximizes expected cumulative reward, allowing effective learning in noisy environments.
@@ VII-C Results and Discussion
TABLE III: Real-world Policy Results ($\mu\pm\sigma$ Per Loop)
Algorithm Steps to Goal $\downarrow$ Safety-human Safety-corridor
Raises $\downarrow$ Raises $\downarrow$
@@
Table III shows DR-MPC outperforming benchmark models in all key metrics—fewer safety raises and faster navigation. The training process (Figure 6) begins with performance nearly identical to the heuristic policy (as expected). During exploration, DR-MPC performs worse but eventually coalesces its experience to achieve superior results. We notice $\beta$ (orange) trends toward 0, while the average $\alpha$ (blue) remains low, indicating the model learns when to best select the MPC action. After 4 hours of data, DR-MPC’s performance begins to plateau. While more data would likely add marginal improvements, social navigation is a long-tail problem, and even in simulation with deterministic humans, safety-human raises never reach zero.
Qualitatively, DR-MPC exhibits key crowd-navigation behaviors: smoothly weaving around still humans (Figure 8 left) and deviating from its path and slowing down for moving humans (Figure 1 and Figure 8 right). In human-free areas, DR-MPC switches to MPC path tracking.
@@
VII-D Limitations and Future Work
One limitation is the manual tuning of the heuristically defined reward function, particularly balancing path-tracking and human-avoidance rewards. For instance, overemphasizing collision penalties impedes progress along the path; Imitation Learning methods bypass this limitation by replicating expert behaviour [37]. Future work includes incorporating reward learning methods, such as Inverse RL [38] and preference learning [11], to enable DR-MPC to acquire more human-aligned behaviors.
Another limitation involves false collision detections due to modeling humans as circles. False positives occur because humans are often wider shoulder-to-shoulder than front-to-back. False negatives arise when extended feet during walking or resting postures cause the center point to misrepresent their state. Future work includes skeleton detection to better capture nuances, improving collision accuracy.
Lastly, we acknowledge that our ‘real-world’ experiments capture only a subset of human behaviors expected in true ‘in-the-wild’ scenarios. Our results reflect interactions with people familiar with the robot. Behaviors such as stopping to observe the robot or intentionally obstructing its path are not encompassed in our current environment. In future work, we aim to conduct ‘in-the-wild’ experiments to evaluate the robot’s performance with more diverse interactions.
@@ Appendix A Human Avoidance Network
For each human trajectory $\mathbf{q}^{t-H_{i}:t}_{i}=\begin{bmatrix}\mathbf{p}^{t-H_{i}}_{i}\dots\mathbf{p}^{t}_{i}\end{bmatrix}$, we sequentially process it from time $t-H_{i}$ to $t$ using a GRU to generate an embedded trajectory $\mathbf{e}^{i}_{\text{traj}}$. This step handles human trajectories of varying lengths, embedding them into a uniform latent space. Next, with $\mathbf{E}_{\text{traj}}=\begin{bmatrix}\mathbf{e}^{1}_{\text{traj}}\dots\mathbf{e}^{n_{t}}_{\text{traj}}\end{bmatrix}$, we use three MLPs to generate the queries $\mathbf{Q}_{\text{traj}}$, keys $\mathbf{K}_{\text{traj}}$, and values $\mathbf{V}_{\text{traj}}$. We then apply multi-head attention using the scaled dot-product attention: $\text{MultiHead}(\mathbf{Q}_{\text{traj}},\mathbf{K}_{\text{traj}},\mathbf{V}_{\text{traj}})$. The output of this module is $\mathbf{E}_{\text{HH}}$, a $n_{t}\times d_{HH}$ tensor, where $d_{HH}$ is the dimension of the human-human embeddings.